You can afford a $900 per month mortgage payment. You've found a 30 year loan at 6% interest.
a) How big of a loan can you afford?
$

b) How much total money will you pay the loan company?
$

c) How much of that money is interest?

Answer by Theo(13342)

6% compounded monthly = 6/12 = .5% per month.
30 years * 12 = 360 months.
present value of 900 paid at the end of each month is equal to 150,112.45.
that's how big a loan you can afford.
the total payments would be 360 * 900 = 324,000.
total interest paid on the loan would be 324,000 minus 150,112.45 = 173,887.55.
you can use the online calculator at https://arachnoid.com/finance/ to get the same answer.
here are the results from using that calculator.

the calculator tells you the present value.
you then have to multiply the payments by 360 to get the total payments.
you then have to subtract the present value to get the total interest.